OpenStudy (anonymous):

solve the system of linear equations in two unknowns : 2x+5y=-1 3x-2y=3

OpenStudy (msmr):

I would multiply the first equation by 3 and the bottom equation by 2 so that it looks like this: 3(2x + 5y = -1) --> 6x + 15y = -6 2( 3x - 2y = 3) --> 6x - 4y = 6 And then subtract the equations: 19y =-12 y = - 12/19 and then substitute that into one of the other equations and solve for x